# A low-rank non-convex norm method for multiview graph clustering

### v1

@@ 1 Introduction
Multiview data clustering is a difficult machine learning task that requires combining data from several data sources or perspectives to cluster data points precisely. In many disciplines, including computer vision, neurobiology, and social network research, multi-view data are frequently encountered. The underlying relationships between viewpoints may not be captured well by conventional clustering approaches that examine each view independently, yielding less-than-ideal clustering outcomes. Thus, it’s crucial to create processes that can efficiently utilize information from various points of view.
Another method that has been extensively utilized for multi-view data analysis is tensor decomposition [1]. The underlying relationships between views can be captured via latent components that can be extracted using tensor decomposition to represent multi-view data as high-order tensors. Popular tensor decomposition techniques, such as non-negative matrices and tensor factorization (NMF and NTF), have been used in multi-view clustering [2].
In this paper, we propose a new method for multi-view clustering that uses a non-convex tensor norm. The proposed method (CGMVC-NC) utilizes the tensor structure of multi-view data and introduces a non-convex tensor norm to capture the correlations between views. The non-convexity of the tensor norm allows it to model complex and non-linear relationships between views. Compared to traditional methods, the proposed method achieves better clustering accuracy on several benchmark datasets.
The contributions of this paper are summarized as follows:
@@
    We propose a novel multi-view clustering method that performs spectral embedding and tensor representation with a parametric non-convex tensor norm.
@@ 2 Preliminaries
The transpose of $\mathcal{X}$ denoted by $\mathcal{X}^{T}\in\mathbb{R}^{N_{2}\times N_{1}\times N_{3}}$ is obtained by transposing each frontal slice and then reversing the order of transposed frontal slices 2 through $N_{3}$.
@@
$\mathcal{X}^{T}*\mathcal{X}=\mathcal{X}*\mathcal{X}^{T}=\mathcal{I}.$
Where *** referred here [3] to the T-product. A tensor is called, f-diagonal if each of its frontal slices is a diagonal matrix. The T-SVD decomposition theorem [3, 4, 5] states that $\mathcal{X}$ can be factorized as:
@@
where $*\left\|.\right\|_{*}$ is the nuclear norm of matrix, which is exactly the sum of the singular values. It is proven [6] to be a valid norm and the tightest convex relaxation to the norm of the tensor average rank within the unit ball of the tensor spectral norm.
@@
where, .$\|.\|_{\gamma}$ represents the Gamma quasi-norm for matrices [8] , for $Z\in\mathbb{R}^{N_{1}\times N_{2}}$
@@ 3 Related work
The multi-view clustering [9] approaches can be roughly divided into two types based on how the similarity graph is constructed: subspace segmentation-based methods and graph-based methods.
@@
Two exemplary of single view clustering techniques are low rank subspace clustering (LRR) [10] and sparse subspace clustering (SSC) [11]. Several multi-view subspace clustering techniques [12, 13] have been proposed based on LRR and SSC. i.e, according to the experimental results in [14], the LRR miscategorizes different data points than SSC. Thus, low-rank and sparsity criteria must be coupled to capture both the global and local structure of the data [14]. In [14], the authors proposed a multi-view subspace clustering method based on the coupled low-rank and sparse representation (CLRS) model.
Treating data as matrices rather than as a full tensor implies that the relationship between subspaces is overlooked; thus, various approaches [15, 16, 17] transform data into a third-order tensor to capture high-order view correlations.
Multi-view graph clustering aims to find a fusion graph of all views, then applies the classical clustering methods on the fusion graph, as spectral clustering or graph-cut algorithms to produces the final clustering. In [18], canonical correlation analysis (CCA) is used to identify the common underlying structure between different views. Although it does not give importance to the weight, in [19], it adds new hyper parameters. In AMGL method [20], there are no hyper parameters, it can learn the weight by itself. In [21, 22], these methods are based on the generalization of the Laplacian rank constrained to give a similarity matrix with exactly the desired number of components. As in the subspace based method [23, 24, 25, 26], data is transformed into a tensor to capture the high-order view correlations within the data.
Noise and redundant information are typically mixed with the original characteristics in practice. Consequently, the learned consensus similarity graph may be incorrect. To overcome this problem, [26], proposed an approach called Consensus graph learning (CGL), which trains an adaptive neighbor graph in a new low-dimensional embedding space rather than the original feature space. First, it learns the view-specific similarity graph $S^{(v)}$ using adaptive neighbor graph learning:
$\forall i=1,...,n\quad$ $\min_{S^{(v)}_{i}}\sum_{j=1}^{n}\left\|\mathbf{x}_{i}^{(v)}-\mathbf{x}_{j}^{(v)}\right\|_{2}^{2}S^{(v)}_{ij}+\gamma s_{ij}$ (3)
$s^{(v)}_{ij}\geq 0,\mathbf{1}^{T}\mathbf{s}^{(v)}_{i}=1,j=1,\ldots,n,$
@@
$\min_{F^{(v)},\mathcal{T}}$ $-\lambda\sum_{v=1}^{V}\operatorname{tr}\left(F^{(v)}A^{(v)}F^{(v)^{T}}\right)+\frac{1}{2}\|\mathcal{F}-\mathcal{T}\|_{F}^{2}+\|\mathcal{T}\|_{w,*}$ (4)
s.t. $F^{(v)^{T}}F^{(v)}=I_{c},v=1,\ldots,V$
$\mathcal{F}=[\overline{F}^{(1)}\overline{F}^{(1)^{T}};\cdots;\overline{F}^{(n)}\overline{F}^{(n)^{T}}],$
@@
$s_{ij}\geq 0,\mathbf{1}^{T}\mathbf{s}_{i}=1,i,j=1,\ldots,n.$
@@ 4.2 Graph construction
  • Cosine similarity: $s_{ij}^{(v)}=\frac{\mathbf{x}_{i}^{(v)^{T}}\mathbf{x}_{j}^{(v)}}{\left\|\mathbf{x}_{i}^{(v)}\right\|_{2}\left\|\mathbf{x}_{j}^{(v)}\right\|_{2}}$
@@ 4.3 Problem formulation
$\min_{F^{(v)},\mathcal{T}}$ $\left[-\lambda\sum_{v=1}^{V}\operatorname{tr}\left(F^{(v)}A^{(v)}F^{(v)^{T}}\right)+\frac{1}{2}\|\mathcal{F}-\mathcal{T}\|_{F}^{2}+\rho\|\mathcal{T}\|_{t-\gamma}\right]$ (7)
[…]
$\mathcal{F}=[\overline{F}^{(1)}\overline{F}^{(1)^{T}};\cdots;\overline{F}^{(V)}\overline{F}^{(V)^{T}}].$
@@ 4.4 Solution
where $\eta_{i}$ is the Lagrangian multiplier of the Lagrangian problem of (3) associated to the constraint $\mathbf{1}^{T}\mathbf{s}_{i}=1$.
@@
[…]
where $H^{(v)}=\lambda A^{(v)}+\frac{1}{2}P^{(v)}\left(\mathcal{T}(:,v,:)+\mathcal{T}(:,v,:)^{T}\right)P^{(v)}-\frac{1}{2}P^{(v)}\overline{F}^{(v)}\overline{F}^{(v)\top}P^{(v)}$.
Remark: The problem is now an optimization problem of trace on a manifold. The matrix $H^{(v)}$ depends on the optimization problem $F^{(v)}$, which makes finding the optimal solution of $F^{(v)}$ thought to solve. Another difficulty, is that the function with respect to $F^{(v)}$ is not smooth (which is due the the inverse of the norm and hence, non differentiability on point zero), and thus not smooth. To manage to give an approximation, we use the Picard method, and since this is a sub-problem, we propose to consider $H^{(v)}$ as a function of the previous iteration, by this consideration, and since $H^{(v)}$ is symmetric, the solution would be well known, according to the Ky Fan’s Theorem [28], are the eigenvectors normalized corresponding to the $c$ largest eigenvalues of $H^{(v)}$.
@@
$\min_{s_{i}}||s_{i}+\frac{1}{2V\gamma}d_{i}||_{2}^{2}$ (19)
$\mathbf{1}^{T}\mathbf{s}_{i}=1,\;s_{ij}\geq 0,\;j=1,\ldots,n,$
@@ 4.5 Algorithm
1:Initialize $S$ by Adaptive neighbor without the rank constraint.
@@
3:     Update $Q$triangleright Using Ky Fan’s Theorem
4:     update $S$triangleright Using Eq. (20)
@@
5:     Update $F^{(v)}$triangleright Solution of Eq. (14)
6:     Update $P^{(v)}$ and $\overline{F}^{(v)}$triangleright Using Eq. (12) and (13)
7:     Compute $\mathcal{F}=[\overline{F}^{(1)}\overline{F}^{(1)^{T}};\cdots;\overline{F}^{(n)}\overline{F}^{(n)^{T}}]$
8:     update $\mathcal{T}$triangleright Using Eq. (16)
@@
10:Compute $S$triangleright Via Algorithm 1
@@ 5.5 Parameter setting
In our suggested model, the parameters are $\gamma$ is chosen from the range $[10^{-4},10^{-3},10^{-2},\\ 10^{-1},1,10,10^{2}]$ , we found that setting $\gamma$ to $0.1$ works best in all cases. For $\lambda$, we follow the suggested in the CGL paper and we set it to $\frac{1}{\sqrt{min(N_{1},N_{3})}}$. For the parameter $\rho$, we’ve tried in the$[10^{-4},10^{-3},10^{-2},10^{-1},1,10,10^{2}]$, the experiments suggests that the value has not a decisive effect.
@@ 5.6 Comparison result and analysis
Our proposed approach performs well on the benchmark datasets. It outperforms the other methods on the accuracy evaluation metric, and does generally well, on the rest of the evaluation metrics. It performs better the the CGL on almost all the metrics over the different dataset.

### v2

@@ 1 Introduction
Multiview data clustering is a challenging machine learning task that requires integration data from several data sources or perspectives to accurately group data points. Multiview data naturally arise in various domains, including computer vision, neurobiology, and social network research. The underlying relationships between viewpoints may not be captured well by conventional clustering approaches that examine each view independently, yielding less-than-ideal clustering outcomes. Consequently, developing techniques that effectively leverage information across multiple views is crucial.
One widely used approach for multiview data analysis is tensor decomposition[1]. By representing multiview data as high-order tensors, tensor decomposition extracts latent components that capture the relationships between views. Popular techniques, such as non-negative matrix factorization (NMF) and non-negative tensor factorization (NTF), have been successfully applied to multiview clustering [2].
In this paper, we propose a new method for multiview clustering that utilizes a non-convex tensor norm. The proposed method CGMVC-NC utilizes the tensor structure of multiview data and introduces a non-convex tensor norm to capture the correlations between views. The non-convexity of the tensor norm allows it to model complex and non-linear relationships between views. Compared to traditional methods, the proposed method achieves high clustering accuracy on several benchmark datasets.
The contributions are summarized as follows:
@@
    We propose a novel multiview clustering method that performs spectral embedding and tensor representation with a parametric non-convex tensor norm.
@@ 2 Preliminaries
The transpose of $\mathcal{X}$ denoted by $\mathcal{X}^{\top}\in\mathbb{R}^{N_{2}\times N_{1}\times N_{3}}$ is obtained by transposing each frontal slice and then reversing the order of transposed frontal slices 2 through $N_{3}$.
@@
$\mathcal{X}^{\top}*\mathcal{X}=\mathcal{X}*\mathcal{X}^{\top}=\mathcal{I}.$
Where * referred here [3] to the t-product. A tensor is called, f-diagonal if each of its frontal slices is a diagonal matrix. The T-SVD decomposition theorem [3, 4, 5] states that $\mathcal{X}$ can be factorized as:
@@
where $\left\|.\right\|_{*}$ is the nuclear norm of matrix, which is exactly the sum of the singular values. It is proven [6] to be a valid norm and the tightest convex relaxation to the norm of the tensor average rank within the unit ball of the tensor spectral norm.
@@
where, $\|.\|_{\gamma}$ represents the Gamma quasi-norm for matrices [8] , for $Z\in\mathbb{R}^{N_{1}\times N_{2}}$
@@ 3 Related work
Multiview clustering [9] approaches can be roughly divided into two types based on how the similarity graph is constructed: subspace segmentation-based methods and graph-based methods.
@@
Low-Rank Representation LRR [10] and Sparse Subspace Clustering SSC [11] are two widely used single-view subspace clustering techniques. Building on these methods, several multiview subspace clustering approaches have been developed [12, 13]. Among them, the Coupled Low-Rank and Sparse Representation CLRS model, proposed in [14], integrates both low-rank and sparsity constraints. Experimental results in [14] show that LRR and SSC misidentify different data points, highlighting their complementary nature. Therefore, coupling low-rank and sparsity criteria is crucial for effectively capturing both the global and local structures of the data.
A different approach [31], considers regarding graphs as Markov chain, with a co-supervising strategy with structure information. Treating data as matrices rather than as a full tensor implies that the relationship between subspaces is overlooked; thus, various approaches [15, 16, 17] transform data into a third-order tensor to capture high-order view correlations.
Multi-view graph clustering aims to find a fusion graph of all views, then applies the classical clustering methods on the fusion graph, as spectral clustering or graph-cut algorithms to produces the final clustering. In [18], canonical correlation analysis CCA is used to identify the common underlying structure between different views. Although it does not give importance to the weight. [19] adds new hyper parameters. [20] presented a method named AMGL without introducing any hyper parameters. AMGL can learn the weight by itself. In [21, 22], these methods are based on the generalization of the Laplacian rank constrained to give a similarity matrix with exactly the desired number of components. As in the subspace based method [23, 24, 25, 32], data is transformed into a tensor to capture the high-order view correlations within the data.
Noise and redundant information are typically mixed with the original characteristics in practice. Consequently, the learned consensus similarity graph may be incorrect. To overcome this problem, [32], proposed an approach called Consensus graph learning CGL, which trains an adaptive neighbor graph in a new low-dimensional embedding space rather than the original feature space.
CGL method composed of three steps: First, it learns the view-specific similarity graph $S^{(v)}$ using an adaptive neighbor graph learning:
$\forall i=1,.,n$ $\min_{S^{(v)}_{i}}\sum_{j=1}^{n}\left\|\mathbf{x}_{i}^{(v)}-\mathbf{x}_{j}^{(v)}\right\|_{2}^{2}S^{(v)}_{ij}+\gamma s_{ij}$ (3)
$s^{(v)}_{ij}\geq 0,\mathbf{1}^{\top}\mathbf{s}^{(v)}_{i}=1,j=1,\ldots,n,$
@@
$\min_{F^{(v)},\mathcal{T}}$ $-\lambda\sum_{v=1}^{V}\operatorname{tr}\left(F^{(v)}A^{(v)}F^{(v)^{\top}}\right)+\frac{1}{2}\|\mathcal{F}-\mathcal{T}\|_{F}^{2}+\|\mathcal{T}\|_{w,*}$ (4)
s.t. $F^{(v)^{\top}}F^{(v)}=I_{c},v=1,\ldots,V$
$\mathcal{F}=[\overline{F}^{(1)}\overline{F}^{(1)^{\top}};\cdots;\overline{F}^{(n)}\overline{F}^{(n)^{\top}}],$
@@
$s_{ij}\geq 0,\mathbf{1}^{\top}\mathbf{s}_{i}=1,i,j=1,\ldots,n.$
@@ 4.2 Graph construction
  • Cosine similarity: $s_{ij}^{(v)}=\frac{\mathbf{x}_{i}^{(v)^{\top}}\mathbf{x}_{j}^{(v)}}{\left\|\mathbf{x}_{i}^{(v)}\right\|_{2}\left\|\mathbf{x}_{j}^{(v)}\right\|_{2}}$
@@ 4.3 Problem formulation
$\min_{F^{(v)},\mathcal{T}}$ $\left[-\lambda\sum_{v=1}^{V}\operatorname{tr}\left(F^{(v)}A^{(v)}F^{(v)^{\top}}\right)+\frac{1}{2}\|\mathcal{F}-\mathcal{T}\|_{F}^{2}+\rho\|\mathcal{T}\|_{t-\gamma}\right]$ (7)
$\text{ s.t. }F^{(v)}\in V_{c}(\mathbb{R}^{n}),v=1,\cdots,V$
$\mathcal{F}=[\overline{F}^{(1)}\overline{F}^{(1)^{\top}};\cdots;\overline{F}^{(V)}\overline{F}^{(V)^{\top}}].$
@@ 4.4 Solution
where $\eta_{i}$ is the Lagrangian multiplier of the Lagrangian problem of (3) associated to the constraint $\mathbf{1}^{\top}\mathbf{s}_{i}=1$.
@@
$\text{ s.t. }F^{(v)\top}F^{(v)}=I_{c},$
where $H^{(v)}=\lambda A^{(v)}+\frac{1}{2}P^{(v)}\left(\mathcal{T}(:,v,:)+\mathcal{T}(:,v,:)^{\top}\right)P^{(v)}-\frac{1}{2}P^{(v)}\overline{F}^{(v)}\overline{F}^{(v)\top}P^{(v)}$.
Remark: The problem is now an optimization problem of trace on a manifold. The matrix $H^{(v)}$ depends on the optimization problem $F^{(v)}$, which makes finding the optimal solution of $F^{(v)}$ thought to solve. Another difficulty, is that the function with respect to $F^{(v)}$ is not smooth (which is due the inverse of the norm and hence, non differentiability on point zero), and thus not smooth. To manage to give an approximation, we use the Picard method, and since this is a sub-problem, we propose to consider $H^{(v)}$ as a function of the previous iteration, by this consideration, and since $H^{(v)}$ is symmetric, the solution would be well known, according to the Ky Fan’s Theorem [28], are the eigenvectors normalized corresponding to the $c$ largest eigenvalues of $H^{(v)}$.
@@
[…]
$\mathbf{1}^{\top}\mathbf{s}_{i}=1,\;s_{ij}\geq 0,\;j=1,\ldots,n,$
@@ 4.5 Algorithm
1:Initialize $S$ by an adaptive neighbor without the rank constraint.
@@
3:  Update $Q$. $\triangleright$ Using Ky Fan’s Theorem
4:  update $S$. $\triangleright$ Using Eq. (20)
@@
5:  Update $F^{(v)}$ $\triangleright$ Solution of Eq. (14).
6:  Update $P^{(v)}$ and $\overline{F}^{(v)}$ $\triangleright$ Using Eq. (12) and (13)
7:  Compute $\mathcal{F}=[\overline{F}^{(1)}\overline{F}^{(1)^{\top}};\cdots;\overline{F}^{(n)}\overline{F}^{(n)^{\top}}]$.
8:  update $\mathcal{T}$. $\triangleright$ Using Eq. (16)
@@
10:Compute $S$ $\triangleright$ Via Algorithm 1
@@ 5.5 Parameter setting
In our suggested model, the parameters are $\gamma$ is chosen from the range $10^{-4},10^{-3},10^{-2},\\ 10^{-1},1,10,10^{2}]$ , we found that setting $\gamma$ to $0.1$ works best in all cases. For $\lambda$, we follow the suggested in the CGL paper and we set it to $\frac{1}{\sqrt{min(N_{1},N_{3})}}$. For the parameter $\rho$, we’ve tried in $10^{-4},10^{-3},10^{-2},10^{-1},1,10,10^{2}], the$ experiments suggests that the value has not a decisive effect.
@@ 5.6 Comparison result and analysis
Our proposed approach performs well on the benchmark datasets. It outperforms the other methods on the accuracy evaluation metric, and does generally well, on the rest of the evaluation metrics. It performs better the CGL on almost all the metrics over the different dataset.